• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name pzhu

    Log entry time 01:37:21 on March10,2012

    Entry number 365199

    keyword=start optics data

    since 1H05A harp is not working, we start to take optics run. james pulled out liqued helium already. right now we moved target to carbon and target position to the center of target.